OpenStudy (anonymous):

PLEASE HELP AND WILL GIVE MEDAL! 1.Provide a counterexample that shows when the statement is false. Show all work to verify that the counterexample really makes the statement false. For all integers a, b, and c, if a = b, then .

OpenStudy (nurali):

Counterexample: a=3, b=3, c=0 Proof: 1) Assume if a=b, then a/c = b/c. 2) Let a=3, b=3, c=0. 3) a=b since 3=3 4) a/c = b/c 5) Multiply both sides by c: (ac)/c = b 6) Substitute: [(3)(0)]/c = 3 7) 0/c = 3 8) 0/0 = 3 So, assuming a=b -> a/c = b/c leads to 0/0 = 3, which is false.

OpenStudy (anonymous):

whoa, thank you so much!

OpenStudy (nurali):

Anytime.